OpenStudy (anonymous):

y varies jointly with p and q. When p = 7 and q = 2, y = 70. Find y when p = 8 and q = 4. y = 120 y = 160 y = 200 y = 240

OpenStudy (jhannybean):

jointly means directly i'm guessing?

OpenStudy (lopezking1):

y = k*p*q since it is varying jointly 70 = 2*7*k k = 5 y = 5*8*4 = 160 B

OpenStudy (anonymous):

y varies inversely as three times x. When x = 6, y = 2. Find x when y = 12. x = 1 x = 3 x = 6 x = 9

OpenStudy (anonymous):

@Jhannybean yes, jointly means directly, i think they use jointly when there are two variables

OpenStudy (jhannybean):

okay.

OpenStudy (whpalmer4):

classic examples of joint variation from physics would be force due to gravity or electrostatic attraction between two bodies

OpenStudy (lopezking1):

x=1 The equation of variation then is this: y = 36/3x y = 12/x Above we canceled a factor of 3 in the numerator and denominator. Now we can find x when y = 12: y = 12/x x = 12/y x = 12/12 x = 1. So when y = 12, x = 1.

OpenStudy (jhannybean):

Oh,makes more sense @whpalmer4

OpenStudy (anonymous):

Solve –4x + 7 = x + 2 x = –1 x = 1 x = –3 x = 3

OpenStudy (whpalmer4):

@jhannybean and as a bonus, at no extra charge, you get inverse variation with the square of the distance — call now, operators are standing by to take your order! :-)

OpenStudy (whpalmer4):

@pamelasaucedo you shouldn't post unrelated questions in the same thread, open a new one.

OpenStudy (jhannybean):

@lopezking1 walk her through the steps and let her figure out the answer.

OpenStudy (lopezking1):

Move the - 4x over so it is 7 = 4x + x + 2

OpenStudy (lopezking1):

Then take the 2 over so it is 7 - 2 = 4x + x